the initial velocity of a particle is u (-5/3) m/s² if the acceleration of the object is equal to (3/-4) m/s² and it moves at 3s find the final velocity
1
Expert's answer
2020-12-06T17:26:37-0500

The velocity "v" of an object moving with a constant acceleration at any moment of time "t" is given by the following formula:


"v = u+at"

where "u = -\dfrac{5}{3}m/s" is the initial velocity of the object and "a = -\dfrac{3}{4}m/s^2" is its acceleration. Thus after the "t = 3s", obtain:


"v = -\dfrac{5}{3}-\dfrac{3}{4}\cdot 3 = -\dfrac{47}{12}m/s"

Answer. (-47/12) m/s.